divide (p cubed + 64) divided by (p + 4)
is the answer p squared + 16 - 4p

Answer by venugopalramana(3286)

divide (p cubed + 64) divided by (p + 4)
is the answer p squared + 16 - 4p
USE FORMULA A^3+B^3=(A+B)(A^2-AB+B^2)
P^3+4^3=(P+4)(P^2-4*P+4^2)
HENCE WE GET
(P^3+64)/(P+4)=(P+4)(P^-4P+16)/(P+4)=P^2-4P+16